### v2

# IKUN for WMT24 General MT Task: LLMs Are here for Multilingual Machine Translation

Abstract
This paper introduces two multilingual systems, IKUN and IKUN-C, developed for the general machine translation task in WMT24. IKUN and IKUN-C represent an open system and a constrained system, respectively, built on Llama-3-8b and Mistral-7B-v0.3. Both systems are designed to handle all 11 language directions using a single model. According to automatic evaluation metrics, IKUN-C achieved 6 first-place and 3 second-place finishes among all constrained systems, while IKUN secured 1 first-place and 2 second-place finishes across both open and constrained systems. These encouraging results suggest that large language models (LLMs) are nearing the level of proficiency required for effective multilingual machine translation. The systems are based on a two-stage approach: first, continuous pre-training on monolingual data in 10 languages, followed by fine-tuning on high-quality parallel data for 11 language directions. The primary difference between IKUN and IKUN-C lies in their monolingual pre-training strategy. IKUN-C is pre-trained using constrained monolingual data, whereas IKUN leverages monolingual data from the OSCAR dataset. In the second phase, both systems are fine-tuned on parallel data sourced from NTREX, Flores, and WMT16-23 for all 11 language pairs.
1 Introduction
Large language models (LLMs) Touvron et al. (2023); Dubey et al. (2024); Jiang et al. (2023); OpenAI (2023) serve as a crucial foundation for a wide range of applications. One significant advantage of LLMs is their ability to be applied across various tasks, thereby simplifying deployment processes. However, the application of LLMs to multilingual machine translation (MT) presents several challenges:
  • Most LLMs are pre-trained on one or a few dominant languages, making direct fine-tuning on multilingual data insufficient for ensuring optimal performance, particularly for low-resource languages, which are often underrepresented in the training data.
  • It remains unclear whether these LLMs, primarily pre-trained on a limited number of languages, effectively facilitate transfer learning across different languages Tan et al. (2024).
  • The large-scale nature of most LLMs presents significant challenges for efficient fine-tuning, particularly for researchers and practitioners with limited computational resources.
In the WMT24 general MT task Kocmi et al. (2024), our objective is to assess the capability of LLMs for multilingual MT, as an alternative to training bilingual systems from scratch Wu et al. (2023). This paper provides a detailed account of how we developed our final multilingual system using LLMs for both the constrained and open tracks.
Firstly, we identified that certain LLMs exhibit inefficiencies in tokenizing sentences from languages that are underrepresented in the pre-training data. To address this, we extended the existing vocabulary to reduce the tokenized sentence length, thereby enhancing training efficiency. Secondly, we enriched the LLMs with knowledge across the 10 target languages through continued pre-training. This step is particularly crucial for underrepresented languages, as it facilitates transfer learning. Finally, we fine-tuned the models using high-quality parallel datasets across all 11 pairs.
Through this streamlined approach, our constrained multilingual system, IKUN-C, secured 6 first-place and 3 second-place rankings in the automatic evaluation. Our open multilingual system, IKUN, achieved 1 first-place and 2 second-place rankings across the open and constrained tracks. These encouraging results demonstrate that LLMs can be effectively adapted for multilingual MT, broadening access to speakers of diverse languages.
Figure 1: Tokenizer efficiency for various LLMs and languages. The larger the length ratio is, the less efficient the tokenizer is. We add new sub-words, from Chinese, Japanese, Hindi and Icelandic, to the Mistral-v0.3 vocabulary to construct the IKUN-C vocabulary. IKUN uses the Llama-3 tokenizer without any modification.
2 Pre-trained LLM
LLMs are pre-trained on extensive web-scale data, encompassing a vast repository of general knowledge applicable to various tasks. Previous studies Xu et al. (2024a, b) have demonstrated that LLMs can substantially enhance the performance of multilingual MT. Building on this insight, we adopt a pre-trained LLM as the foundation for our system.
IKUN is an open system developed with meticulous consideration of available resources and system capabilities. For this purpose, we selected Llama-3 Dubey et al. (2024), one of the most advanced open-source LLMs available at the time of this competition. Due to constraints on computational resources, we opted for the 8B version [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B instead of the 70B version. A significant factor in our choice of Llama-3 was its strong support for multilingual applications, as evidenced by the efficiency with which its tokenizer handles all 11 languages involved in this competition (See §3). We also tried the instruct version, but it is worse than the pre-trained version.
IKUN-C is a constrained system based on Mistral-7B-v0.3 Jiang et al. (2023), one of the three LLMs permitted for the constrained track. Prior to selecting Mistral-7B-v0.3222https://huggingface.co/mistralai/Mistral-7B-v0.3, we conducted continuous pre-training on all three allowed LLMs — namely, Llama-2-7B, Llama-2-13B Touvron et al. (2023), and Mistral-7B — using a subset of our monolingual data (approximately 1B tokens). The pre-training loss demonstrated that Mistral-7B outperformed Llama-2-7B and performed comparably to Llama-2-13B, leading us to select it as our architecture of choice.
3 Tokenizer Efficiency
A significant challenge in applying LLMs to multilingual MT lies in the efficiency of their tokenizers. These models are typically pre-trained on one or a few dominant languages, and when their tokenizers are applied to low-resource languages, they produce disproportionately long sequences of sub-words. This inefficiency leads to excessive GPU memory consumption during training.
To evaluate the efficiency of the tokenizer, we focus on comparing the length differences between tokenized English sentences and their corresponding non-English counterparts. Specifically, we define the length ratio as:
$\mathrm{length}\ \ \mathrm{ratio}=\frac{\mathrm{len(tokenizer(}x\mathrm{))}}{\mathrm{len(tokenizer(}y\mathrm{))}}$
where $y$ represents the English sentence, and $x$ denotes the paired non-English sentence. A smaller length ratio (close to 1) is desired, since it means that the tokenizer can encode the non-English sentence as efficient as the English sentence.
To facilitate a comparison of length ratios across different languages, English-centric multilingual data is essential. Fortunately, the FLoRes-200 dataset Costa-jussà et al. (2022) possesses this characteristic. In the devtest and test sets of FLoRes-200, every English sentence is paired with translations in various other languages. We concatenate all sentences from the devtest set and compute the length ratio for each language, as illustrated in Figure 1. We also include NLLB’s tokenizer Costa-jussà et al. (2022) for a comparison.
We can observe: (1) NLLB consistently exhibits the smallest length ratio across all languages, likely due to its extensive optimization for hundreds of languages, thus serving as a lower bound in this context. (2) Mistral-v0.3 and Llama-3 demonstrate a notably high length ratio for Hindi, suggesting that Hindi is underrepresented in the pre-training data. (3) Compared to NLLB, the tokenizer of Mistral-v0.3 is significantly less efficient for Chinese, Japanese, Hindi, and Icelandic.
We opted to expand the vocabulary by incorporating new sub-words to reduce the length of tokenized sentences, thereby enhancing training efficiency. However, this approach introduces a trade-off between the addition of new sub-words and training performance. The embeddings for the newly introduced sub-words are initially untrained, and a substantial increase in sub-words may necessitate additional iterations of continuous pre-training. Consequently, our strategy for adding sub-words prioritizes those from languages with higher length ratios.
For our open system, IKUN, we didn’t modify its tokenizer, since Llama-3 tokenizer is already efficient enough, only except for Hindi. For our constrained system, IKUN-C, we expanded its vocabulary for Chinese, Japanese, Hindi, and Icelandic through the following steps: (1) Generate a new vocabulary of 12K sub-words using monolingual data from the past two years for these four languages from News Crawl [^3] https://data.statmt.org/news-crawl/; and (2) Merge the new vocabulary with the original. The efficiency of the resulted IKUN-C tokenizer is shown in Figure 1, demonstrating more efficiency for these four languages, especially for Hindi.
Language pair Num. Language pair Num.
cs-uk 8768 uk-cs 8768
ja-zh 12858 zh-ja 12858
en-zh 36647 zh-en 34650
en-cs 30120 cs-en 28123
en-de 35564 de-en 33567
en-hi 11020 hi-en 9023
en-is 8010 is-en 6013
en-ja 18113 ja-en 16116
en-ru 32840 ru-en 30843
en-es 13808 es-en 11811
en-uk 11961 uk-en 9964
en-fr 4006 fr-en 4006
Total: 429457
Table 1: Number of parallel sentences.
Hyper-parameter Continuous pre-training Finetuning
sampling probability cs,de,en,es,hi,is,ja,ru,uk,zh =
0.1,0.13,0.1,0.13,0.08,0.05,0.08,0.13,0.08,0.12
duration 60K steps 1 epoch
batch size 64 128
sequence length 2048 max source length=512, max target length=512
learning rate (lr) 2e-5 2e-4
warmup ratio 0 0.01
weight decay 0.01 0.01
lr scheduler cosine inverse_sqrt
training type full finetuning LoRA $r=64$ for all layers
Table 2: Experimental setting for continuous pre-training and subsequent finetuning.
System Metric cs-uk en-cs en-de en-es en-hi en-is en-ja en-ru en-uk en-zh ja-zh
MetricX $\downarrow$ 2.4 4.3 2.0 3.5 7.1 4.9 4.3 4.7 4.7 4.2 6.2
CometKiwi $\uparrow$ 0.648 0.618 0.641 0.666 0.499 0.657 0.669 0.649 0.622 0.624 0.519
IKUN-C AutoRank $\downarrow$ 3.0 4.7 3.8 3.4 5.5 3.7 3.9 3.9 3.9 3.5 5.5
Place in constrained $\downarrow$ 2 5 1 1 1 1 3 1 1 2 2
MetricX $\downarrow$ 1.6 3.7 1.8 3.3 9.4 4.3 3.7 4.1 3.7 4.0 5.4
CometKiwi $\uparrow$ 0.664 0.638 0.668 0.687 0.428 0.666 0.696 0.675 0.661 0.646 0.544
IKUN AutoRank $\downarrow$ 2.3 3.9 3.0 2.8 7.7 3.2 3.1 3.2 2.8 3.1 4.4
Place in constrained&open $\downarrow$ 2 5 4 3 5 1 6 3 2 5 6
[…]
4 Experiments
We mainly follow the pipeline from Xu et al. (2024a), i.e. continuous pre-training on monolingual data for all 10 languages, and followed by fine-tuning on parallel data for all 11 language pairs.
4.1 Continuous Pre-training
Given that our selected LLMs, specifically Llama-3-8B and Mistral-7B-v0.3, are primarily pre-trained on English, it is necessary to incorporate knowledge from other languages through further pre-training, with particular emphasis on low-resource languages. Additionally, the word embeddings for the newly introduced sub-words in IKUN-C must also undergo training.
For the open system IKUN, we utilize monolingual data from the Oscar dataset Suárez et al. (2020), covering all 10 target languages. We adopt the sampling strategy outlined by Xu et al. (2024a), described as:
$P(l)\propto(\frac{D_{l}}{\sum_{l^{\prime}\in L}D_{l^{\prime}}})^{\frac{1}{T}}\quad s.t.\quad\sum_{l^{\prime}\in L}P(l^{\prime})=\frac{9}{10}$
where $D_{l}$ represents the number of words in language $l$ [^4] https://huggingface.co/datasets/oscar-corpus/OSCAR-2301, $T$ is the temperature parameter (set to 6), and $L$ denotes the set of all languages except English. The sampling probability for English is fixed at 1/10. The experimental settings for continuous pre-training are detailed in Table 2. We approximately pre-trained IKUN on an additional 8B tokens.
In the constrained system, only the provided data sources are permitted for use [^5] https://www2.statmt.org/wmt24/mtdata/. The IKUN-C system utilizes monolingual data from the News Crawl dataset for 9 languages, with the exception of Spanish, as the use of Spanish monolingual data from News Crawl is restricted. For Spanish, we incorporate monolingual data from the Leipzig Corpora Goldhahn et al. (2012). Additionally, for Hindi, we augment the dataset with monolingual data from News Commentary due to the relatively limited amount of Hindi data available in the News Crawl dataset. This adjustment is crucial because Hindi is underrepresented in the pre-training of Mistral-7B-v0.3. Our experimental settings closely align with those detailed in Table 2.
4.2 Subsequent Fine-tuning
Previous studies Wu et al. (2024); Zhou et al. (2023) have demonstrated that the quality of fine-tuning data is a critical factor in achieving optimal performance. Liao et al. (2021) further indicates that increasing the amount of back-translation data does not necessarily lead to better outcomes. In light of this, we exclusively utilize high-quality parallel data for the fine-tuning phase.
The high-quality parallel data is primarily sourced from FloRes-200 Costa-jussà et al. (2022), NTREX-128 Federmann et al. (2022), and previous WMT16-23 general MT/news tasks Kocmi et al. (2023, 2022); Akhbardeh et al. (2021); Barrault et al. (2020, 2019, 2018); Bojar et al. (2017, 2016).
[…]
Past WMTs: Additionally, we extract parallel sentences from the development and test sets of the WMT16-23 general MT tasks, provided they contain the necessary translation directions. For these sentences, we employ a bidirectional fine-tuning strategy.
The statistics for all parallel sentences are presented in Table 1. Notably, all systems are fine-tuned at the sentence level. Given that WMT development and test sets are at the document level, models could alternatively be fine-tuned at the document level or reformatted into a conversational structure for fine-tuning. This latter approach might be more effective for context-aware translation, as WMT24 applies context-based human evaluations. We reserve this exploration for future work. The fine-tuning setting is listed in Table 2.
4.3 Results
We present the preliminary results reported by Kocmi et al. (2024) in Table 3, which includes four evaluation metrics. Both MetricX-23-XL Juraska et al. (2023) and CometKiwi-DA-XL Rei et al. (2023) have demonstrated strong correlations with human evaluation Freitag et al. (2023). AutoRank Kocmi et al. (2024), a normalized composite metric derived from MetricX and CometKiwi, scales the scores of each metric linearly to span the range from 1 to the total number of systems in a given language pair. The final automatic ranking is obtained by averaging these normalized scores. AutoRank can thus be considered a measure of the overall rank across all systems and tracks. Additionally, we report the rankings of our systems across various tracks.
It is noteworthy that both of our systems are multilingual, designed to handle all language pairs. The IKUN-C system, in particular, demonstrates promising performance in the constrained track, achieving 6 first-place and 3 second-place finishes. In both the open and constrained tracks, IKUN maintains strong performance, securing 1 first-place and 2 second-place positions, even when compared to systems that may leverage additional open-source data or specialize in a limited set of language pairs.
5 Conclusion
In this paper, we present a methodology for effectively adapting pre-trained LLMs to the task of multilingual machine translation. Our approach involves three primary steps: (1) expanding the vocabulary to accommodate languages that are underrepresented in the pre-training data, when necessary; (2) continuing pre-training the LLM on monolingual data to enhance its knowledge of underrepresented languages and to train the embeddings of newly introduced sub-words; and (3) fine-tuning the LLM on high-quality parallel data. Our experimental results demonstrate the efficacy of this straightforward pipeline, with IKUN-C securing 6 first-place finishes in the constrained track, and IKUN achieving 1 first-place ranking in both the open and constrained tracks.
Acknowledgments
We thank eBay Inc. for the computation support. This research was funded in part by the Netherlands Organization for Scientific Research (NWO) under project number VI.C.192.080.